Right now im trying 2 work with finding:
The Value of X,then classify the triangle by its angle measures
Equation: 72,5x,x

Answer by tutorcecilia(2152)

The sum of the angles of a triangle = 180
Angle #1 =72
Angle #2=5x
and Angle #3=x
.
So, (72)+(5x)+ (x)=180 [solve for x]
72+6x=180
6x=108
x=18
.
Plug (x=18) back into the value of each angle and solve:
Angle #1 =72
Angle #2=5x =5(19)=90
and Angle #3=x=18
.
72+90+18=180
180=180 [checks out]